OpenStudy (anonymous):

Integral X^4 sinX dX

OpenStudy (solomonzelman):

use integration by parts. Integral f d g = f g Integral g d f f= x^4 dg= sin(x) dx df = 4x^3 dx g = -cos(x)

OpenStudy (solomonzelman):

Go for it...